3x-2y = -1 X tay=2 find the valuels of a that makes the system has ④ unique solution. ® No solution
Answer #1

-- =) 3x-2y = -1 (1) x +2y = 2 (2) For Ist equation 3x-2y 2y = 3at1 =y= 3 x + 1/ :y=mxtc m=3₂ ; ; / For and equation xray = 2So, da € Real number and ²-0-667 no (B) For solution these two lines must be parallel and do not overlap with each other i.e.